### v1

# Growth of Structural Lengthscale in Kob Andersen Binary Mixtures: Role of medium range order

Abstract
A central and extensively debated question in glass physics concerns whether a single, growing lengthscale fundamentally controls glassy dynamics, particularly in systems lacking obvious structural motifs or medium range crystalline order (MRCO). In this work, we investigate structural and dynamical lengthscales in supercooled liquids using the Kob Andersen binary Lennard Jones (KALJ) model in two compositions: 80:20 and 60:40. We compute the dynamical lengthscale from displacement–displacement correlation functions and observe a consistent growth as temperature decreases. To explore the static counterpart, we use a structural order parameter (SOP) based on the mean field caging potential. While this SOP is known to predict short time dynamics effectively, its bare correlation function reveals minimal spatial growth. Motivated by recent findings that long time dynamics reflect collective rearrangements, we perform spatial coarse-graining of the SOP and identify an optimal lengthscale $L_{max}$ that maximises structure–dynamics correlation. We show that the structural correlation length derived from SOP coarse-grained over $L_{max}$ exhibits clear growth with cooling and closely tracks the dynamical lengthscale, especially for A particles in the 80:20 mixture and for both A and B particles in the 60:40 system. Our results reconcile the previously observed absence of static length growth in MRCO-free models like KALJ by highlighting the necessity of intermediate range structural descriptors. Furthermore, we find that the particles with larger structural length growth also correspond to species with latent crystallisation tendencies, suggesting a possible link between structural order, dynamics, and incipient crystallisation.
I Introduction
Supercooled liquids exhibit a dramatic slowdown in dynamics upon cooling and ultimately form an amorphous solid, or glass Debenedetti2001 ; Cavagna2009 ; Berthier2011 ; Ediger2000 . Despite sustained progress, the microscopic mechanisms driving this glass transition remain debated. A central challenge is to rationalise the rapid increase in the dynamics, spanning many orders of magnitude, with a corresponding static structural signature tanakamrco_kawasaki_prl_2007 ; Coslovich2007 ; royall2008 ; tanakamrco_shintani_natphys_2008 ; Karmakar2009 ; Tanaka2010 ; keys2011 ; Coslovich2011 ; Sastry2014 ; bapst2020 ; Manoj_prl_2021 ; Sharma2022 . This issue has motivated extensive efforts to identify growing structural and dynamical correlation lengths, a key theme in modern glass physics shintani2006 ; tanakamrco_kawasaki_prl_2007 ; tanakamrco_critical_natmat_2010 ; Dunleavy2012 ; MalinsLifetimes2013 ; russo_tanaka ; Tarjus2016pre ; Tah_PRL_2018 .
Dynamical lengthscales are relatively well established, often quantified through multi-point correlation functions such as the four-point dynamic structure factor Karmakar2010 ; Berthier2011 ; Tah_dis_dis ; Paula2023 or the displacement–displacement correlation function  Donati1999PRL ; Tah_dis_dis ; Paula2023 . In contrast, static lengthscales remain elusive, largely because conventional pair correlation functions like radial distribution function ($g(r)$) exhibit little change upon supercooling.
Over the past two decades, considerable effort has been devoted to identifying static lengthscales that grow alongside the well established dynamical correlation length in glass forming systems. Tanaka and coworkers have extensively studied the growth of static lengthscale in systems exhibiting medium range crystalline order (MRCO), such as those with Steinhardt bond orientational order parameterSteinhardt1983 in 3D and hexatic order in 2D and found that these systems show a stronger coupling between static and dynamic lengthscales tanakamrco_shintani_natphys_2008 ; tanakamrco_kawasaki_prl_2007 ; tanakamrco_kawasaki_JPCM_2010 ; tanakamrco_critical_natmat_2010 ; tanakamrco_leocmach_natcomm_2012 ; tanakamrco_boo_jncs_2012 ; tanakamrco_hu_pree_2016 . Note that although in 2D systems, the presence of MRCO is quite common and easy to detect, it is often tricky to find the same in 3D systems. Advanced techniques for measuring static lengthscales, such as the point-to-set (PTS) method, were designed to be order-agnostic biroli_thermodynamic_2008 ; Hocky2012 ; Biroli2013 ; Berthier2016 ; Tarjus2016pre . Tah et al. Tah_PRL_2018 investigated the PTS length in glassy systems, distinguishing between those with and without medium range crystalline order (MRCO). Their work revealed a sublinear scaling of the static PTS length with the dynamical correlation length for systems lacking MRCO, while for MRCO systems, these two lengths were found to grow commensurately. Royall and co-workers have shown that for polydisperse systems in deep supercooling, the static lengthscale connected to the defective icosahedra order grows modestly Hallett2018 .
However, for MRCO-free systems like the 3D Kob Andersen binary Lennard Jones mixture, this correspondence often breaks down. Early studies by Marinari and Pitard Marinari2005 and La Nave et al. LaNave2006 linked cooperative relaxation to persistent clusters and inherent structure landscapes, respectively. Coslovich and Pastore used Voronoi tessellation to identify locally preferred structures (LFS) which are connected to the slow dynamics and showed that their cluster size grows with decreasing temperature Coslovich2007 . Royall and coworkers used their proposed Topological cluster classification (TCC) Malins2013JCP to identify LFS connected to slow particles and studied the growth of these clusters Malins2013JCP ; Dunleavy2012 ; MalinsLifetimes2013 ; Royall2014 ; Crowther2015 ; Hallett2018 ; royall2008 . Karmakar et al. KARMAKAR20121001 showed that the minimum eigenvalue of the hessian matrix $\lambda_{min}$ plays a crucial role in defining and understanding the static length at low temperature where an-harmonic effect is absent. Chakrabarty et al. Chakrabarty2017 introduced a new structure agnostic methodology called block analysis to study the static and dynamic lengthscales. Nevertheless, all these studies show that in systems without medium range crystalline order (MRCO) these structural lengthscales grow more slowly than their dynamical counterparts.
In the KALJ model, the topological cluster classification (TCC) studies identified the 11A bicapped square-antiprism cluster, which is assumed to be a precursor of the CsCl structure, as an LFS associated with slow dynamics, but its spatial extent does not scale with the dynamical correlation length Malins2013JCP ; MalinsLifetimes2013 ; Royall2014 . More recent investigations suggest that in the KALJ model, the particles do not have a CsCl formation tendency between the two species, but the A-type particles show a tendency to form a face-centred cubic (fcc) structure Ujjwal_jcp_2016 . However, these motifs are not easily captured by standard bond-orientational or TCC-based descriptors. This disconnect raises an important question: does the lack of observed static lengthscale growth stem from a true absence of underlying order, or does it reflect the limitations of existing local structural metrics that fail to capture intermediate-range correlations?
In this work, we revisit the Kob Andersen binary Lennard-Jones mixture, examining both the canonical 80:20 composition and a modified 60:40 variant. We begin by analysing the structural lengthscale using a recently developed structural order parameter (SOP) that has been shown to correlate well with particle dynamics Manoj_prl_2021 ; Sharma2022 . While this local bare SOP effectively captures short time dynamical behaviour, we find that it does not show a significant growth in spatial correlations upon supercooling and thus fails to reveal a growing static lengthscale. It is important to note that dynamical correlation lengths are typically extracted from long time, not short time, dynamics. Previous studies on both 2D and 3D polydisperse systems by Tong and Tanaka have demonstrated that for their generalised angular structural order, coarse-graining the order parameter in space provides better correlation with long time dynamics. They also showed that the optimal coarse-graining length, at which the structure-dynamics correlation is maximised, increases with decreasing temperature TongTanaka2018 ; TongTanaka2019 ; TongTanaka2020 ; Tanaka2025 . Study involving elastically collective nonlinear Langevin equation theory has shown that the inter relationship between alpha relaxation time and the local cage and collective elastic barriers and also other thermodynamic quantities, is present only when medium range order in the static length is taken into consideration in the theoryken_sch . The study further showed that the static length grows with density. Similarly, investigations using local volume MeiWang2022 , local caging potential mohit_wca , and order parameters derived via unsupervised machine learning QiuArun2025 have shown that meaningful correlations with long time dynamics emerge only after spatial coarse-graining of the static order parameter. Here, we demonstrate that the SOP, when coarse-grained over an optimal lengthscale, exhibits strong correlations with long time dynamics. Notably, the static lengthscale extracted from this coarse-grained SOP grows with decreasing temperature in a manner consistent with the dynamical lengthscale. This growth, for the first time is observed in the MRCO-free KALJ model, provides new evidence that medium range structural correlations, not purely local order, are essential to understanding the dynamical slowdown in glass-forming systems lacking conventional crystalline precursors.
The organisation of the rest of the paper is as follows. Section II provides the simulation details. In Section III, we describe the methodology used in this work. Section IV presents the results. Section V presents a concluding discussion. Two appendices are included for additional details.
II Simulation Details
We simulate the standard 80:20 Kob-Andersen binary Lennard-Jones (KALJ) mixture Kob1994 in three dimensions, consisting of particles of types A and B. The interaction between any pair of particles of types $\alpha$ and $\beta$, separated by a distance $r$, is given by
$U_{\alpha\beta}(r)=4\epsilon_{\alpha\beta}\left[\left(\frac{\sigma_{\alpha\beta}}{r}\right)^{12}-\left(\frac{\sigma_{\alpha\beta}}{r}\right)^{6}\right]$ (1)
where $\alpha,\beta\in\{\text{A},\text{B}\}$ denote the types of the interacting particles. The interaction parameters are set as follows: $\epsilon_{\mathrm{AA}}=1.0$, $\epsilon_{\mathrm{AB}}=1.5$, $\epsilon_{\mathrm{BB}}=0.5$, $\sigma_{\mathrm{AA}}=1.0$, $\sigma_{\mathrm{AB}}=0.8$, and $\sigma_{\mathrm{BB}}=0.88$. All quantities are expressed in reduced LJ units. The number density is fixed at $\rho=1.2$ for the 80:20 mixture and $\rho=1.41$ for the 60:40 mixture. Interactions are truncated and quadratically shifted at a cutoff distance of $r_{c}=2.5\sigma_{\mathrm{AB}}$ to ensure continuity of both the potential and its first derivative Stoddard1973 . Simulations are performed in the NVT ensemble using a Nosé-Hoover thermostat with a time step $\Delta t=0.005$. To minimize finite-size effects, we set the system size to $N=16{,}000$ for structural quantity calculations and $N=500{,}000$ for evaluating the four-point correlation function, following the approach outlined in Reference Karmakar2010 .
III Methodology
III.1 Overlap function and four-point structure factor
The time-dependent overlap function is defined by,
$Q(t)=\frac{1}{N}\sum_{i=1}^{N}\omega(|\mathbf{r}_{i}(t)-\mathbf{r}_{i}(0)|)$ (2)
where the function $\omega(x)$ takes the value 1 if $0\leq x\leq a$ and 0 otherwise and $\mathbf{r}_{i}(t)$ is the position of particle $i$ at time $t$. $a=0.3\sigma$ is typically chosen to account for small particle shifts due to low-amplitude vibrations, which are treated as negligible. This value is comparable to the mean squared displacement (MSD) observed in the plateau region between the ballistic and diffusive regimes. The $\alpha$-relaxation time, $\tau_{\alpha}$, is defined as the time at which $Q(t)$ decays to $1/e$, i.e., $Q(\tau_{\alpha})=1/e$.
The, four-point structure factor $S_{4}({\bf{q}},t)$, which characterises the spatial correlations of the dynamical fluctuations in the wavevector plane q is given by Karmakar2010
$S_{4}(\mathbf{q},t)=\frac{1}{N}\Bigl{[}\langle Q(\mathbf{q},t)Q(-\mathbf{q},t)\rangle-\langle Q(\mathbf{q},t)\rangle^{2}\Bigr{]}$ (3)
Here $Q(\mathbf{q},t)$ is the Fourier Transform of the overlap function given in Eq.2 and can be written as,
$Q(\mathbf{q},t)=\sum_{i=1}^{N}\omega(|\mathbf{r}_{i}(t)-\mathbf{r}_{i}(0)|)e^{i\mathbf{q}\cdot\mathbf{r}_{i}(0)}$ (4)
To obtain the dynamic correlation length we fit the $S_{4}(\mathbf{q},t)$ at $t=\tau_{\alpha}$ to an Ornstein–Zernike form in the small $q$ limit as,
$S_{4}(\mathbf{q},\tau_{\alpha})=\frac{S_{4}(q\!\to\!0,\tau_{\alpha})}{1+(q\xi_{4})^{2}}$ (5)
where $\xi_{4}$ is the dynamic correlation length that grows on approaching the glass transition Karmakar2010 ; Paula2023 . For data analysis, we plot $1/S_{4}(q,\tau_{\alpha})$ versus $q^{2}$; from Eq.5 this should be linear, so an ordinary least-squares fit reliably extracts both the intercept $1/S_{4}(q\!\to\!0,\tau_{\alpha})$ and the slope $\xi_{4}^{2}/S_{4}(q\!\to\!0,\tau_{\alpha})$, which can be then used to obtain $\xi_{4}$.
III.2 Displacement–displacement correlation function
To characterise the spatial extent of correlated motion in supercooled liquids, we use the displacement–displacement correlation function $g_{uu}(r,\tau_{\alpha})$, as studied earlier, Poole1998PhysicaA ; Donati1999PRL ; Paula2023 ; Hallett2018 .
It is defined as:
$g_{uu}(r,\tau_{\alpha})=\frac{\left\langle\sum_{\begin{subarray}{c}i,j=1\\ j\neq i\end{subarray}}^{N}u_{i}(\tau_{\alpha})u_{j}(\tau_{\alpha})\delta\left(r-|\mathbf{r}_{ij}(t)|\right)\right\rangle}{4\pi r^{2}\Delta rN\rho\langle u(\tau_{\alpha})\rangle^{2}}$ (6)
where $u_{i}(\tau_{\alpha})=|\mathbf{r}_{i}(t+\tau_{\alpha})-\mathbf{r}_{i}(t)|$ is the scalar displacement of particle $i$ over time $\tau_{\alpha}$.
To isolate the dynamic part from static density correlations, the computation of the excess correlation function is done by using
$\Gamma_{uu}(r,\tau_{\alpha}))=\frac{g_{uu}(r,\tau_{\alpha})}{g(r)}-1$ (7)
where g(r) is the radial distribution function and is defined as
$g(r)=\frac{\left\langle\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N}\delta\left(r-|\mathbf{r}_{ij}(t)|\right)\right\rangle}{4\pi r^{2}\Delta r\rho N}$ (8)
where $\rho$ is the number density and N is the number of particles. An increasing $\Gamma_{uu}(r,\tau_{\alpha})$ with decreasing temperature indicates growing dynamical heterogeneity near the glass transition. Assuming that $\Gamma_{uu}(r,\tau_{\alpha})$ decays exponentially with distance, then the associated lengthscale can be extracted by integrating $\Gamma_{uu}(r,\tau_{\alpha})$ over space as shown in the equation belowPoole1998PhysicaA ,
[…]
yielding $\xi_{D}$, the characteristic dynamical correlation length.
III.3 Mean field caging potential as structural order parameter
To characterize the static structure of glass forming systems, we introduce the structural order parameter (SOP)Manoj_prl_2021 ; Sharma2022 . The SOP quantifies the extent to which particles are confined by the mean field caging potential resulting from interactions with neighbouring particles, based on the Ramakrishnan–Yussouff free-energy functional RamakrishnanYussouff1979 . Here, we distinguish between the macroscopic SOP $\beta\Phi$ and the microscopic SOP $\beta\phi$ by this notation.
The macroscopic effective caging potential at zero displacement is computed as:
$\beta\Phi=-\rho\int\textbf{dr}\sum_{u,v}C_{uv}(r)\,x_{u}x_{v}\,g_{uv}(r),$ (10)
where $\rho$ represents the number density, and $x_{u}$, $x_{v}$ indicate mole fractions of particle species $u$ and $v$. The partial radial distribution function, $g_{uv}(r)$, is given byHansenMcDonald2013 ,
$g_{uv}(r)=\frac{1}{4\pi r^{2}\rho_{v}N_{u}}\left\langle\sum_{i=1}^{N_{u}}\sum_{j=1}^{N_{v}}\delta(r-|\mathbf{r}_{i}^{u}-\mathbf{r}_{j}^{v}|)\right\rangle$ (11)
where $\rho_{v}$ denotes the species-specific number density, $N_{u}$ and $N_{v}$ represent the total numbers of particles for species $u$ and $v$, respectively, and the angular brackets represent an ensemble average.
The direct correlation function $C_{uv}(r)$ in this study is approximated using the Hypernetted Chain (HNC) closure relation HansenMcDonald2013 :
$C_{uv}(r)=-\beta U_{uv}(r)+(g_{uv}(r)-1)-\ln g_{uv}(r)$ (12)
where $U_{uv}(r)$ denotes the interaction potential between particles of species $u$ and $v$, and $\beta=1/(k_{B}T)$ is the inverse thermal energy.
At the microscopic scale, the effective caging potential experienced by particle $i$ of species $u$ is given as:
$\beta\phi^{i}_{u}=-\rho\int\textbf{dr}\sum_{v}x_{v}C_{uv}^{i}(r)\,g_{uv}^{i}(r)$ (13)
where $\rho$ is the number density, $C_{uv}^{i}(r)$ is the particle-specific direct correlation function, and $g_{uv}^{i}(r)$ represents the local radial distribution function around particle $i$.
To obtain $g_{uv}^{i}(r)$, a Gaussian smoothing method is applied PiaggiValssonParrinello2017 :
$g_{uv}^{i}(r)=\frac{1}{4\pi\rho r^{2}}\sum_{j}\frac{1}{\sqrt{2\pi\delta^{2}}}\exp\left(-\frac{(r-r_{ij})^{2}}{2\delta^{2}}\right)$ (14)
where $r_{ij}$ is the separation between particles $i$ and $j$, and $\delta=0.09\sigma_{AA}$ is the width of the Gaussian smoothing kernel. This approach yields a smooth, noise-reduced local radial distribution function compared to discrete binning.
To avoid unphysical artifacts at short distances in microscopic calculations, we consistently use an approximate expression for the direct correlation function as previously detailed in references mohit_wca ; PatelSharmaBhattacharyya2023 . Specifically, the simplified form adopted is given by
$C_{uv}^{\text{approx}}(r)=g_{uv}(r)-1$ (15)
which maintains computational consistency and physical realism throughout the analysis.
III.4 Structural lengthscale calculation
To quantify the structural lengthscale, we use the microscopic caging potential described before as the structural order parameter, $\beta\phi$. Its structural correlation function is defined as follows
$\bigl{\langle}\beta\phi(0)\,\beta\phi(r)\bigr{\rangle}=\frac{\Bigl{\langle}\displaystyle\sum_{i,j}\beta\phi_{i}\,\beta\phi_{j}\,\delta\!\bigl{(}r-\lvert\mathbf{r}_{ij}\rvert\bigr{)}\Bigr{\rangle}}{4\pi r^{2}\,\Delta r\,\rho\,Ng(r)}$ (16)
The corresponding normalized correlation function isCubuk2017 ; LandauSchuttler1999
$\Gamma_{bare}(r)=\frac{\bigl{\langle}\beta\phi(0)\,\beta\phi(r)\bigr{\rangle}-\bigl{\langle}\beta\phi\bigr{\rangle}^{2}}{\bigl{\langle}(\beta\phi)^{2}\bigr{\rangle}-\bigl{\langle}\beta\phi\bigr{\rangle}^{2}}$ (17)
Integrating $\Gamma_{bare}(r)$ over $r$ yields the characteristic structural lengthscale $\xi_{bare}$,
[…]
which is analogous to the procedure used to obtain the dynamical lengthscale (Eq.9).
III.5 Isoconfigurational ensemble
The iso-configurational ensemble, originally introduced by Harrowell et al. harrowell , consists of multiple trajectories initiated from the same initial particle configuration but with randomised initial momenta drawn from the Maxwell-Boltzmann distribution at the given temperature. This approach eliminates the trivial variations in particle displacements that result from different initial momenta. Any significant differences observed in the trajectory-averaged dynamics can, therefore, be attributed to the properties of the initial configuration rather than to stochastic variations.
[…]
$\mu^{j}(t)=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}\sqrt{(r_{i}^{j}(t)-r_{i}^{j}(0))^{2}}$ (19)
Here, $\mu^{j}(t)$ represents the mobility of the $j^{th}$ particle at time $\tau_{\alpha}$, while $N_{IC}$ denotes the number of independent trajectories. We then quantify the relationship between the initial structure and the dynamics over time using the Spearman rank correlation.
III.6 Spearman rank correlation
Spearman rank correlation, a non-parametric measure of statistical dependence, is computed for a dataset containing $m$ values as follows:
$C_{R}(X,Y)=1-\frac{6\sum d_{i}^{2}}{m(m^{2}-1)}$ (20)
where the squared rank difference is defined as $d_{i}^{2}=R(X_{i})-R(Y_{i})$, with $R(X_{i})$ and $R(Y_{i})$ representing the respective ranks of the raw data points $X_{i}$ and $Y_{i}$.
IV Results
Figure 1: Inverse of the four-point structure factor, $1/S_{4}(\mathbf{q},\tau_{\alpha})$, (Eq.3) plotted as a function of the squared wavevector, $q^{2}$, at various temperatures. Solid lines are Ornstein–Zernike (OZ) fits. (b) Excess displacement-displacement correlation, $\Gamma_{uu}(r,\Delta t)$, (Eq.7) at different temperatures. (c) Excess correlation of bare SOP, $\Gamma_{\text{bare}}(r)$, (Eq.17) at different temperatures. (d) Temperature dependence of three characteristic lengthscales: the dynamical length from OZ analysis of $S_{4}(\mathbf{q},\tau_{\alpha})$, $\xi_{4}$ (Eq.5); the dynamical lengthscale from $\Gamma_{uu}(r,\tau_{\alpha})$, $\xi_{D}$ (Eq.9); and the bare static lengthcale from $\Gamma_{bare}(r)$, $\xi_{bare}$ (Eq.18). For comparison, the lengthscales are all scaled. The scaling factor for $\xi_{4}$ is 4.28, $\xi_{D}$ is 0.24, & $\xi_{bare}$ is 0.71. With a decrease in temperature, both dynamical lengthscales exhibit similar growth, while the bare static lengthscale shows weaker growth. The colour coding in (b) & (c) matches that of (a).
Figure 2: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu)$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu$, calculated at times scaled by the $\alpha$-relaxation time, $t/\tau_{\alpha}$, for A-type particles at different coarse-graining lengths, $L$. (a) Result at high temperature, $T=0.70$ and (b) at low temperature, $T=0.45$. The colour coding in (b) matches that of (a).
Figure 3: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu(\tau_{\alpha})$, calculated at $\alpha$-relaxation time as a function of coarse-graining length, $L$, (a) for A-type particles and (b) for B-type particles at different temperatures. The colour coding in (b) matches that of (a) (c) Temperature dependence of $L_{max}$, the value of $L$ in (a) and (b) at which $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$ peaks, shows strong growth for A particles and weaker growth for B particles. Lines are power‐law fits.
Figure 4: Results for 80:20 KALJ system: Top panels: Snapshots at $T=0.45$ showing structural and dynamical fields for A-type particles: (a) The bare SOP, $\beta\phi$, appears spatially noisy, lacking visible large-scale structure. (b) Coarse-graining the SOP, $\overline{\beta\phi}$, with $L_{max}$, shows the emergence of patches of high and low SOP values. (c) The particle mobility field, $\mu(\tau_{\alpha})$, calculated using Eq.19 over the $\alpha$-relaxation time shows low mobility domains that spatially mostly overlap with high SOP regions in (b), indicating a structure–dynamics link. Bottom Panels: Excess correlation functions corresponding to the fields in top panels at different temperatures. (d) The excess correlation of the bare SOP, $\Gamma_{\text{bare}}(r)$, (Eq.17) confirms the absence of significant growth in bare static lengthscale. (e) The excess correlation of the coarse-grained SOP, $\Gamma_{\text{CG}}(r)$, (Eq.22) shows an increasing correlation length with decreasing temperature, consistent with the emergence of extended structural patches. (f) The excess displacement–displacement correlation function, $\Gamma_{uu}(r,t_{\alpha})$, (Eq.7) exhibits growth with decreasing temperature. The colour coding in (e) and (f) matches that of (d)
Figure 5: Temperature dependence of the dynamical lengthscale, $\xi_{D}$, and coarse-grained structural lengthscale, $\xi_{CG}$, in the 80:20 KALJ system. $\xi_{D}$ is obtained from $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.9); and $\xi_{CG}$ obtained from $\Gamma_{CG}(r)$, (Eq.23). For comparison, the lengthscales are all scaled. The scaling factor for $\xi_{D}$ A-type is 0.16, $\xi_{D}$ B-type is 0.15, $\xi_{CG}$ A-type is 0.91, & $\xi_{CG}$ B-type is 7.70.
Figure 6: (a) Distribution of the bare SOP, $\beta\phi$, for all particles (black, circle) and for those belonging to 11A TCC clusters (red, square). The 11A particles exhibit a pronounced shift toward deeper cages. (b) Overlap function (Eq.2) for all particles (black, circle), 11A particles (red, square), and the 2000 most deeply caged particles (top 12.5%) (green, triangle). Both 11A and 2000 deeply caged particles show significantly slower dynamics than the average.
Figure 7: Excess displacement–displacement correlation $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.7) at different temperatures in the 60:40 mixture. (a) A-type particles.(b) B-type particles. The colour coding in (b) matches that of (a).
[…]
Figure 8: Results for 60:40 KALJ system: Top panels: Spearman rank correlation, $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained structural order parameter, $\overline{\beta\phi}$, and mobility, $\mu(\tau_{\alpha})$, calculated at the $\alpha$-relaxation time for various coarse-graining lengths $L$, for (a) A-type particles at various temperatures and (b) B-type particles at various temperatures. (c) Temperature dependence of $L_{max}$, the value of $L$ in (a) and (b) at which $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha})$ peaks, shows strong growth for A particles and B particles. Lines are power law fits. Middle panels: Snapshot at T=0.45 for the structural fields: (d) bare SOP, $\beta\phi$, for A-type particles; (e) coarse-grained SOP, $\overline{\beta\phi}$, for A-type particles; (f) bare SOP, $\beta\phi$, for B-type particles; and (g) coarse-grained SOP, $\overline{\beta\phi}$, for B-type particles. Bare SOP appears spatially noisy, lacking visible large scale structure. Coarse-graining the SOP, with $L_{max}$, shows the emergence of patches of high and low SOP values. Bottom panels: Excess correlation functions corresponding to the fields in the middle panels at different temperatures: (h) bare correlation function, $\Gamma_{\text{bare}}(r)$ (Eq.17), for A-type particles; (i) coarse-grained correlation function, $\Gamma_{\text{CG}}(r)$ (Eq.22), for A-type particles; (j) bare correlation function, $\Gamma_{\text{bare}}(r)$ (Eq.17), for B-type particles; and (k) coarse-grained correlation function, $\Gamma_{\text{CG}}(r)$ (Eq.22), for B-type particles. While bare SOP fields show little spatial correlation due to their noisy nature, the coarse-grained fields exhibit clear growth in spatial correlations, consistent with the emergence of structural patches. The colour coding in (b),(h),(i),(j) and (k) matches that of (a).
Figure 9: Temperature dependence of the dynamical lengthscale, $\xi_{D}$, and coarse-grained structural lengthscale, $\xi_{CG}$, in the 60:40 KALJ system for the A and B types of particles. $\xi_{D}$ is obtained from $\Gamma_{uu}(r,\tau_{\alpha})$, (Eq.9); and $\xi_{CG}$ is obtained from $\Gamma_{CG}(r)$, (Eq.23). For comparison, the lengthscales are all scaled. The scaling factor for $\xi_{D}$ A-type is 1.17, $\xi_{D}$ B-type is 0.28, $\xi_{CG}$ A-type is 0.16, & $\xi_{CG}$ B-type is 0.15.
We first investigate the correlation between the dynamical and static lengthscales in the 80:20 Kob Andersen binary Lennard Jones mixture (KALJ).
There are different formalisms to calculate the dynamical lengthscale. In Fig.1(a) we plot $1/S_{4}(\mathbf{q},\tau_{\alpha})$ (Section III.1) as a function of $q^{2}$ and from this fitting using Ornstein Zernike expression (Eq.5) we obtain the dynamical correlation length $\xi_{4}$. Using the displacement-displacement correlation function, $g_{uu}(r,\tau_{\alpha})$, we calculate the excess displacement-displacement correlation function, $\Gamma_{uu}(r,\tau_{\alpha})$ (Eq.7) plotted in Fig.1(b). Integration of this excess correlation gives us another measure of the dynamic lengthscale, $\xi_{D}$ (Eq.9) Poole1998PhysicaA .
Here, we study the static lengthscale as described by the recently developed structural order parameter, the depth of the mean field caging potential, $\beta\phi$  Manoj_prl_2021 ; Sharma2022 . Like the dynamical lengthscale, in Fig.1(c), we plot the excess correlation of the bare SOP, given by $\Gamma_{bare}(r)$ (Eq.17). The lengthscale $\xi_{bare}$ is then obtained from the area under the curve (Eq.18). In Fig.1(d), we plot the two dynamical lengthscales and the static lengthscale. We find that both the dynamical lengthscales grow similarly; however, the static lengthscale, almost does not grow.
In order to further investigate how the structural and the dynamical lengthscales correlate, we first examine the structure-dynamics correlation. This correlation is studied using isoconfigurational ensembles. Isoconfigurational ensembles (Section III.5) were designed by Harrowell and coworkers harrowell to disentangle the influence of the initial structure from thermal noise in the dynamics. In this part of the analysis, to obtain cleaner statistics, we separately deal with the A and the B particles in the system. The results of the structure-dynamics correlation, given by the Spearman rank correlation, $C_{R}({\beta\phi},\mu)$ between the bare SOP and the mobility for the $A$ particles are given in the main text and that for the $B$ particles are given in Appendix II. As shown in Fig.2(a), $C_{R}({\beta\phi},\mu)$ between the local per particle level bare SOP with the mobility for the $A$ particles is high at shorter times, but at longer times there is a drop in the correlation. Note that we define. This drop is stronger at lower temperatures as seen in Fig.2(b). The results are similar for the $B$ particles (Fig.11 in Appendix II). This indicates that the local bare SOP has strong predictive power for short time dynamics, but its influence diminishes at longer times, particularly at low temperatures.
As mentioned in Section III.1 and Section III.2, the dynamical heterogeneity lengthscale is typically evaluated at the time scale, $\tau_{\alpha}$ where the spatial correlations of particle displacements are most pronounced and dynamic heterogeneity reaches its maximum. Berthier_Biroli_PRE_2005 Thus, any static order parameter which does not show a good correlation at that timescale may not be able to predict the correct static length scale.
Recently, there has been a large number of studies TongTanaka2018 ; das_possible_2018 ; TongTanaka2019 ; TongTanaka2020 ; Tanaka2025 ; MeiWang2022 ; QiuArun2025 , including us mohit_wca , where it is shown that if we coarse-grain the SOP, then the structure-dynamics correlation increases at longer times. The coarse-grained SOP,$\overline{\beta\phi}_{i}(L)$, is defined as
[…]
where the sum over $j$ includes all particles within a distance $L$ of the tagged particle $i$. This coarse-graining reflects the influence of a particle’s local environment, weighted exponentially by spatial proximity, and thus captures collective structural effects beyond the single-particle scale. In Fig.2(a), we plot the Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu)$ against scaled time $t/\tau_{\alpha}$ for different values of coarse-graining length. We find that at short times for higher $L$ values, the correlation is poor, and then at longer times, the correlation increases, and at even longer times, it decreases. We also find that this increase in correlation at longer times initially increases with $L$ and eventually, for higher $L$ values, the correlation drops. This implies there is an optimum value of $L$, we refer to as $L_{max}$ where the correlation is maximum. When we compare the plots in Fig.2(a) and 2(b), we find that $L_{max}$ increases with decreasing temperature. It is evident from Fig.2 that the value of $L_{max}$ also depends on the time at which the mobility is calculated.
As the dynamical lengthscale is calculated at $\tau_{\alpha}$, we next analyse the correlation and the corresponding value of $L_{max}$ for the mobility at $t=\tau_{\alpha}$. In Fig.3(a) and Fig.3(b) we plot Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu(\tau_{\alpha}))$, between the coarse-grained SOP and the mobility calculated at $t=\tau_{\alpha}$ for the $A$ and $B$ particles, respectively, against different coarse-graining length, $L$. This plot clearly shows that at each temperature, there is an optimum value of $L$ for which the correlation is maximum. Both for A and B particles, this $L_{max}$ value increases with a decrease in temperature.
However, at low temperatures, the $L_{max}$ value for the $B$ particles is less than that of the $A$ particles (Fig.3(c)), and this difference increases with a decrease in temperature. Note that although we study the A and the B particles separately, while coarse-graining the SOP, if the neighbouring particle is of a different species, its value is included in the coarse-graining process. Thus, the lower value of $L_{max}$ for the $B$ particles is not due to a lower concentration of the $B$ particles. This just shows that, compared to that for the A particles, fewer particles contribute to the long time dynamics of the $B$ particles.
In Fig.4(a-c), we show the snapshots of the single particle bare SOP, coarse-grained SOP and also the mobility at $t=\tau_{\alpha}$ for the A particles at $T=0.45$. We find that the snapshot of the single particle bare SOP (Fig.4(a)) is highly fluctuating data and has no correlation with the mobility at long times (Fig.4(c)). However, when we plot SOP, which is coarse-grained over $L_{max}=3.56$ we find regions of high and low values of SOP (Fig.4(b)) which visually correlates with the regions of high and low values of mobility (Fig.4(c)) It is this correlation between the coarse-grained SOP and the mobility that is quantified in terms of Spearman rank correlation values given in Fig.3(a). From the above plots, a key observation is that once we introduce a coarse-graining length $L_{max}$, the ability of SOP to predict which particles will move significantly over long times improves. Physically, this indicates that the long time rearrangements depend not just on the SOP value of a single particle but also on the SOP values of its surrounding particles, giving rise to the medium range order. The lengthscale of this medium range order also increases with a decrease in temperature, a concept which is in tune with the growing cooperatively rearranging regions.
Our study also suggests that the static order parameter itself changes with the timescale over which we measure the dynamics. While studying short time dynamics, the static order parameter that correlates with the dynamics should be calculated at a single particle level, as seen in Fig.2(a) (black circles) and in earlier studies involving some of us Sharma2022 . Note that a similar observation was also made for the machine learning studies of the softness parameter by Liu and coworkers liu_nature ; Cubuk2017 . However, while studying properties related to long time dynamics like the dynamical correlation length, the corresponding static order parameter is no longer a single particle order parameter but a coarse-grained quantity, as seen in Fig2(b) (pink triangles). This observation is similar to that made in earlier studies QiuArun2025 ; TongTanaka2019 ; MeiWang2022 ; mohit_wca ; ken_sch .
Thus, to extract the correct static lengthscale which is commensurate with the dynamical lengthscale we reassign the SOP value of each particle with its coarse-grained value and calculate the excess correlation between the coarse-grained SOP as,
$\Gamma_{CG}(r)=\frac{\bigl{\langle}\overline{\beta\phi}(0)\,\overline{\beta\phi}(r)\bigr{\rangle}-\bigl{\langle}\overline{\beta\phi}\bigr{\rangle}^{2}}{\bigl{\langle}(\overline{\beta\phi})^{2}\bigr{\rangle}-\bigl{\langle}\overline{\beta\phi}\bigr{\rangle}^{2}}$ (22)
and the corresponding coarse-grained static lengthscale, $\xi_{CG}$ is given by,
[…]
In Fig.4(d), we now plot the $\Gamma_{bare}(r)$, which is the excess correlation between the bare SOP values of the A particles which are calculated at the single particle level, and find that as a function of temperature, it grows moderately, similar to that observed for all particles in Fig.1(c). However, when we plot the $\Gamma_{CG}(r)$, which is the excess correlation between the coarse-grained SOP values, we find that $\Gamma_{CG}(r)$ for the A particles at lower temperatures survives till longer distances. This clearly shows that the static lengthscale related to the coarse-grained order parameter grows. Note that this growth of excess correlation of the coarse-grained SOP at low temperatures can be visually seen in the form of patches of high and low SOP values in Fig.4(b). In Fig.4(f), we also plot the excess displacement-displacement correlation only for the $A$ particles. The plots for the excess correlation for the bare and coarse-grained SOP values and the excess displacement-displacement correlation for the B particles are plotted in Fig.12 in Appendix II.
The static and dynamic lengthscales obtained for the $A$ and $B$ particles are shown in Fig.5. The static lengthscales obtained from $\Gamma_{CG}$ are also separately calculated for the A and the B particles by coarse-graining their SOP values till their respective $L_{max}$ values given in Fig.3(c). We find that the dynamic lengthscale of both species grows similarly. However, the static lengthscale of only the A particles grows like the dynamical lengthscale, but that of the B particles grows moderately, as shown in Fig.5.
Interestingly, many of the earlier studies, trying to find the static correlation length for an 80:20 KALJ system, worked with LFS, like the 11A structure obtained from TCC Malins2013JCP , which is also the ⟨0,2,8⟩ Voronoï polyhedra Coslovich2011 . This structure was found to be long lived, thus contributing to the slow dynamics Royall2014 ; Coslovich2011 . Here we plot the distribution of the SOP of particles in the 11A structure obtained via TCC Malins2013JCP along with the distribution of the SOP of all the particles in Fig.6(a). We find that the SOP values of the 11A particles are shifted towards a higher caging potential, suggesting that these particles are indeed in a deeper caging potential. However, the distribution is also quite wide, which implies that, unlike proper crystals, the local arrangements of the particles in the 11A structure vary over a wide range. We plot the overlap function of the particles which are in the 11A structure along with the overlap function of all the particles in Fig.6(b). We find that, indeed, the particles belonging to the 11A structure are slower compared to the average dynamics. Note that at $T=0.45$ on average, there are about $12.5\%$ particles in the 11A structure. For a system of 16,000 particles, this comes to about 2000 particles. So we calculate the overlap function for the 2000 hardest particles and plot it in Fig.6(b). We find that they are marginally slower than the 11A particles. Thus, we show that our SOP can also identify slow moving particles. Our study reveals that the failure of the earlier studies in obtaining the growing static lengthscale in the 80:20 KALJ system is due to two possible reasons. The first is that the required coarse-graining to take care of the medium range order was not done, and the second is that many studies focused on LFS structures, which for this system is centred around the B particles that we know do not show much growth.
We next study the growth of the static and dynamic lengthscale in a 60:40 mixture. This system is studied at a higher density $\rho=1.41$ so that the temperature range is similar to that of the 80:20 mixture (Fig.10 in Appendix I). In Fig.7(a) and Fig.7(b), we plot the excess displacement-displacement correlation functions of the A and the B particles, respectively. We find that similar to the 80:20 mixture, both for the A and B particles, the excess correlation grows with the lowering of the temperature. In Fig.8(a) and Fig.8(b), we plot the Spearman rank correlation of the coarse-grained SOP and the mobility at $t=\tau_{\alpha}$ for the A and B particles, respectively. Both show that there is an optimum value of coarse-graining length $L_{max}$ where the correlation is maximum, and $L_{max}$ grows with a decrease in temperature. Unlike in the case for the 80:20 mixture, in this case (Fig.8(c)), the $L_{max}$ for the $B$ particles grows almost as much as that for the $A$ particles.
In Fig.8(d) and Fig.8(f), we show the snapshots of the bare SOP values for the A and the B particles, respectively, at $T=0.45$. These show that the values of the SOP fluctuate at the single particle level. In Fig.8(e) and Fig.8(g), we show snapshots of the coarse-grained SOP for the A and the B particles, respectively. We find that the coarse-grained SOP does show patches of regions with low and high SOP values. In Fig.8(h) and Fig.8(j), we plot $\Gamma_{bare}$ (Eq.17) at different temperatures, and they show a weak temperature dependence. However, $\Gamma_{CG}$ (Eq.22) for the A and the B particles show that with a decrease in temperature, the correlation survives till a longer distance suggesting a growing correlation length (Fig.8(i) and Fig.8(k)). To quantify these results in Fig.9 we plot the dynamic and the static lengthscales for the A and the B particles. We find that in this case, both the static and the dynamical lengthscales for both species grow together.
V Conclusions
In this work, we investigate the relationship between structural and dynamical lengthscales in Kob Andersen binary Lennard Jones (KALJ) mixtures with 80:20 and 60:40 compositions, focusing on how these lengthscales evolve as the systems are supercooled. Using displacement–displacement correlation functions, we compute the dynamic correlation length, $\xi_{D}$, which exhibits clear growth with decreasing temperature, consistent with increasing dynamic heterogeneity near the glass transition. We further find that the dynamical lengthscales for both A and B particles grow in a similar fashion when analysed separately.
To examine the static counterpart, we analyse lengthscales derived from a recently developed structural order parameter (SOP) based on a mean field caging potential Manoj_prl_2021 ; Sharma2022 . When using the local bare SOP at a single particle level, the static lengthscale shows limited growth, echoing earlier observations Hocky2012 ; KARMAKAR20121001 ; Biroli2013 ; Berthier2016 ; Yaida2016 ; MalinsLifetimes2013 ; Royall2014 . This result is initially surprising given previous studies involving some of us, demonstrated a strong correlation between this SOP and dynamics Sharma2022 ; mohit_wca . To resolve this, we analyse the structure–dynamics correlation across different timescales. We find that while the local SOP correlates well with short time dynamics, its correlation weakens over long timescales. This is because long time rearrangements are influenced not only by the SOP of a given particle but also by the SOP of its neighbours, indicating the emergence of medium range order.
By coarse-graining the SOP over a spatial lengthscale, $L$, we identify an optimal coarse-graining length, $L_{max}$, that maximises structure–dynamics correlation at $\tau_{\alpha}$ time scale. We find that this $L_{max}$ grows with a decrease in temperature. For the 80:20 mixture, this growth is more pronounced for the A particles than for the B particles, while for the 60:40 mixture, both species show comparable growth. The corresponding static correlation length, extracted from the spatial correlations of the coarse-grained SOP, also increases significantly at low temperatures. Moreover, we find that for the 80:20 mixture, only the static lengthscale of A particles grows with cooling, and in the 60:40 mixture, the static lengthscales of both species grow in tandem with the dynamical lengthscale. At this point we do not completely understand why the structural lengthscale for the B particles do not grow but as discussed later there might be some correlation with the underlying crystalline order.
These findings provide important insight into the long standing question of whether structural order underlies the slowing down of dynamics in glass forming liquids. Our results indicate that the structural order parameter itself changes depending on the timescale over which the dynamical quantities are studied. For short time dynamics, the bare structural order parameter is a good order parameter; however, to understand the properties connected to the long time dynamics, we need the medium range structural order parameter captured through coarse-grained SOPs. This helps reconcile previous inconsistencies in static lengthscale growth observed in KALJ systems and supports the idea that dynamic slowdown is accompanied by the growth of structural order, albeit not purely local in nature.
Finally, we investigate whether the observed structural correlations are linked to underlying crystalline order. Previous analyses involving some of us showed that in the 80:20 mixture, the A particles exhibit a tendency to form fcc structures, while no evidence of CsCl-like ordering was observed Ujjwal_jcp_2016 . This indicates that B particles do not participate in the crystallization process in this composition. In contrast, the 60:40 mixture displays a tendency to form a CsCl structure, which involves both A and B particles Ujjwal_jcp_2016 . Interestingly, neither composition exhibits locally favored structures that correspond to the underlying crystalline order. Nevertheless, we find that the static length scale grows predominantly for the particle species associated with the dominant crystallization tendency, namely, the $A$ particles in the 80:20 mixture, and both $A$ and $B$ particles in the 60:40 mixture. This suggests a potential link between the growth of structural correlations and latent crystallization pathways. However, further investigations are needed to establish a definitive connection.
Overall, our study establishes a coherent framework that connects structural and dynamical perspectives in glass formation. It highlights the importance of choosing the correct spatial scale when analysing structural indicators relevant to the dynamical quantity. Future work could explore whether alternative structural indicators that have captured static lengthscale growth TongTanaka2019 ; Tanaka2025 in other systems can reveal similar signatures in the KALJ systems. Also, it will be important to test if other structural order parameters like the machine learning softness liu_nature ; Cubuk2017 show growth in correlation length when coarse-grained.
Appendix I: Dynamics in 80:20 and 60:40 composition in KALJ systems
Figure 10: The $\alpha$-relaxation time obtained from the overlap function (Eq.2), $\tau_{\alpha}$, for two systems: 80:20 composition at $\rho=1.2$ and 60:40 composition at $\rho=1.41$ as a function of temperature. The solid lines are guides to the eye.
In our study, we examine two KALJ systems at different compositions: the well-known 80:20 composition and a 60:40 composition whose density is tuned to $\rho=1.41$ so that both systems exhibited quantitatively similar relaxation dynamics in the same temperature range as shown in Fig.10. Although the local structural motifs differ due to compositional variation, the overall relaxation behaviour is comparable in both systems.
Appendix II: Lengthscale studies of B-type particles in 80:20 composition
In Fig.11 we plot the Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu)$ against time scale by $\tau_{\alpha}$ which quantifies the relationship between the coarse-grained SOP and particle mobility for $B$ particles. As shown in Fig.11(a), for $L=0$ (i.e., the bare SOP), the correlation is initially strong at short times but decays rapidly, indicating that the bare SOP captures structural features relevant only to short time dynamics. This decay becomes more pronounced at lower temperatures, as illustrated in Fig.11(b). To predict dynamics over longer timescales, it becomes necessary to incorporate medium range structural information. This is evident from the increased structure–dynamics correlation when the SOP is coarse-grained over a spatial length $L$; up to a certain optimum value, $L_{max}$, the coarse-grained SOP retains a stronger correlation with mobility at longer times. Here, in Fig.12, we show the snapshots of the structural and dynamical parameters and their respective excess correlations for the B particles. We find that even for the coarse-grained SOP, the size of the patches representing high and low SOP values is smaller than that of the A particles (see Fig.4(b)). This leads to the smaller values of $L_{max}$ and $\xi_{CG}$ for the B particles compared to the A particles.
Figure 11: Spearman rank correlation, $(C_{R}(\overline{\beta\Phi},\mu))$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu$, calculated at times scaled by the $\alpha$-relaxation time, $t/\tau_{\alpha}$, for B-type particles at different coarse-graining lengths, $L$. (a) Result at high temperature, $T=0.70$ and (b) at low temperature, $T=0.45$. The colour coding in (b) matches that of (a).
Figure 12: Snapshots of structural and dynamical properties and their corresponding excess correlations for B-type of particles for 80:20 KALJ system at T=0.45. (a) Snapshot of the bare SOP, $\beta\phi$, appears spatially noisy, lacking visible large scale structure. (b)The excess correlation of the bare SOP, $\Gamma_{\text{bare}}(r)$, (Eq.17) confirms the absence of growth in bare static lengthscale.(c) Snapshot of the coarse-grained SOP, $\overline{\beta\phi}$, with $L_{max}$, shows emergence of some small patches of high and low SOP values. However these patches are much smaller than that observed for the A-type of particles (see Fig.4(b)) (d)The excess correlation of the coarse-grained SOP, $\Gamma_{\text{CG}}(r)$, (Eq.22) shows an increasing correlation length with decreasing temperature. Consistent with the size of the patches the static lengthscales obtained from the excess correlation are smaller than that of the A-type of particles (see Fig.5) (e)Snapshot of the particle mobility field, $\mu(\tau_{\alpha})$, calculated using Eq.19 over the $\alpha$-relaxation time shows low and high mobility domains. (f) The excess displacement–displacement correlation function, $\Gamma_{uu}(r,t_{\alpha})$, (Eq.7) exhibits growth with decreasing temperature.The colour coding in (d) and (f) matches with (b).
Acknowledgements.
S. M. B. thanks, Science and Engineering Research Board (SERB, Grant No. SPF/2021/000112 ) and CSIR-NCL, Pune for funding. S. K. thanks DST INSPIRE for Fellowship. M. S. thanks CSIR for fellowship. S. M. B. would like to thank Smarajit Karmakar and Chandan Dasgupta for discussions. S. K. would like to thank Palak Patel and Anoop Mutneja for discussions.